A rectangular storage unit has dimensions of : 1m by 2m by 3m. If each linear dimension is increased by the same amount.&
a) What increase would result in a new storage unit with a volume of 10 times the orginal?&
b) How many possible solutions are there to this problem?&
1
Expert's answer
2010-12-16T14:05:41-0500
Initial volume is V = a*b*c = 1 *2 * 3 = 6 m3. 10V = 60 m3
a) Increace of multiplication of linear dimentions by 10 times would lead to increacing the volume by 10 times; xa*yb*zc / abc = 10.
b) there are infitinite number of solutions of this problem; ( you can choose any combination of x, y, z, the only condition is xyz = 10)